Q: On a journey from planet P to Venus a rocket is traveling at a constant speed during this journey the rocket Travels pass the moon in 4 days The distance from planet period to the moon is 4.8 x 10^5 km.
i) Fimd the distance travelled by rocket in 12 days. giv answer in standard form.
ii) Given that the distance between planet p and Venus is 4.8x10^7,find the time taken , in days for the journey. :)

Answer by CubeyThePenguin(3113)

(4.8 * 10^(5))/4 = 1.2 * 10^(5) km traveled every day

i) 1.2 * 10^(5) * 12
= 2.4 * 10^(6)
= 2,400,000 km in 12 days

ii) (4.8 * 10^(7))/(1.2 * 10^(5)) = 400 days